# Constructing g-computation estimators: two case studies in selection bias

### v1

@@ Abstract
G-computation is a useful estimation method that can be adapted to address various biases in epidemiology. However, these adaptations may not be obvious for some complex causal structures. This challenge is an example of the much wider issue of translating a causal diagram into a novel estimation strategy. To highlight these challenges, we consider two recent cases from the selection bias literature: treatment-induced selection and co-occurrence of biases that lack a joint adjustment set. For each case study, we show how g-computation can be adapted, described how to implement that adaptation, show some general statistical properties, and illustrate the estimator using simulation. To simplify both the theoretical study and practical application of our estimators, we express the proposed g-computation estimators as stacked estimating equations. These examples illustrate how epidemiologists can translate identification results into an estimation strategy and study the theoretical and finite-sample properties of a novel estimator.
@@ 1 Introduction
A practical challenge to empirical epidemiologic research is translating from a causal diagram identification result to a strategy for estimation. While the epidemiologic and statistical literature is filled with novel estimators, these estimators are often tailored to specific scenarios. For example, an estimator may not deal with multiple sources of bias (e.g., confounding, selection bias, missing data), or assume there is an overall sufficient adjustment set for all the sources of systematic errors [1]. The gap between drawing the causal diagram and developing an estimation strategy may dissuade some researchers from considering these novel estimators, even when they are thought to produce more reliable estimates relative to traditional regression methods.
@@
Here, we consider methods to address selection bias as a case study of translating from identification results to estimation. Recent work has focused on clarifying the definition of selection bias [2, 3, 4], and provided graphical rules for addressing selection bias [5, 6]. However, these studies do not immediately provide corresponding estimators to recover treatment effects from selection bias. When faced with selection bias, epidemiologists may default to inverse probability weighting estimators, as they allow for one to independently model different sources of bias (e.g., inverse probability of treatment weights for confounding, inverse probability of censoring or sampling weights for different types of selection bias) [7]. While inverse probability weighting provides a fairly general recipe, there are challenges to its application [8]. First, an epidemiologist must correctly specify each model used to construct inverse probability weights, which can be difficult as it requires sufficient but, oftentimes, distinct background knowledge for each source of bias. Second, the order in which different inverse probability weights are constructed can differ depending on the context [9]. Third, inverse probability weighting estimators are generally less efficient (i.e., have larger variances and wider confidence intervals) relative to competing estimators [10]. These issues can be avoided by adopting outcome-model-based approaches (e.g., g-computation [11]), which only require correct specification of a model for the outcome process (avoiding separate models for each source of bias and concern over the ordering of estimation of different weights) and is generally more efficient [8, 12]. However, g-computation may need to be adapted for certain selection bias mechanisms or under the joint occurrences of biases (e.g., confounding and selection bias) that lack a sufficient adjustment superset [1, 13]. Therefore, adaptations of the standard g-computation algorithm are needed for causal structures under these scenarios.
@@
Using two recent examples in the selection bias literature as case studies, we illustrate how epidemiologists can translate identification results into novel estimators. Here, we focus on g-computation. For completeness, we rederive the motivating identification results, propose estimators based on those identification results, prove some general properties of our estimators, and illustrate their application using Monte Carlo simulation experiments. For the first case study, we consider treatment-induced selection bias. In the second case study, we study the joint occurrence of confounding and selection bias where there is not an overlapping adjustment set. While estimators are developed for these specific cases, the central ideas of this work can be used to build g-computation estimators in other contexts.
@@ 2 Estimating Equations
M-estimators provide four benefits for us. First, M-estimators provide a general recipe for estimating the variance of parameters via the so-called empirical sandwich variance estimator [15, 19]. By jointly solving a stack of estimating equations, the empirical sandwich variance estimator appropriately incorporates the dependence of parameters on other parameters. Essentially, we can estimate the variance of the parameter of interest, while correctly incorporating the uncertainty in estimation of the outcome model for g-computation. This feature allows us to avoid more computationally intensive methods for variance estimation, such as the bootstrap [20, 21]. Note that this feature also applies to inverse probability weighting estimators [22], where the variance of the parameter of interest depends on the uncertainty in estimation of the propensity scores. Second, M-estimation theory can simplify consistency and asymptotic normality proofs for estimators. Hence, we can more easily prove certain statistical guarantees for our estimator as the sample size approaches infinity. Briefly, if the estimating equations are shown to be unbiased at the true $\theta$, it suffices to show that the corresponding estimator is consistent and asymptotically normal following some additional regularity conditions [15]. This underlying statistical theory is what justifies use of the empirical sandwich variance estimator. As we propose novel estimators, we use this second feature to justify use of our estimators. Third, many of the estimators that epidemiologists routinely use can be readily expressed as estimating equations. For example, both the arithmetic mean and generalized linear model can be expressed as estimating equations [8, 15]. Therefore, the underlying pieces of the estimators proposed in this paper are connected to approaches already used by epidemiologists. Finally, there is free and open-source software for the automated computation of M-estimators [23, 24], simplifying their application.
@@ 3 Standard G-computation
Before the case studies, we first review a standard g-computation implementation for the average causal effect with selection bias and how it can be implemented as an M-estimator. Let $Y$ denote the outcome, $A$ denote the treatment, $S$ denote whether a person completed follow-up, and $X$ be a sufficient adjustment set for both confounding and selection bias due to differential loss to follow-up. Note that $Y$ is only observed for those who completed follow-up (i.e., $S=1$). The parameter of interest is $\psi=\mu^{1}-\mu^{0}$, where $\mu^{a}=E[Y^{a}]$, $Y^{a}$ is the potential outcome under treatment $a$, and $E[\cdot]$ is the expected value function. This parameter can be expressed in terms of the observed data as
@@
following assumptions of conditional exchangeability with positivity for both confounding and selection bias, and causal consistency (see Appendix 1 for details) [25, 26, 27].
Using the algorithm described by Snowden et al. adapted for missing outcome [11], we can estimate the average causal effect. First, we fit an outcome model for $Y$ given $A,X$ among those who completed follow-up (i.e., $S=1$). Using that model, we then generate predict values of $Y$ under $a$ for all observations in the study sample (including both $S=1$ and $S=0$), which we refer to as pseudo-outcomes and denote as $\hat{Y}^{a}$. Then we take the mean of the pseudo-outcomes, $\hat{\mu}^{a}=n^{-1}\sum_{i=1}^{n}\hat{Y}^{a}_{i}$, where $n$ is the size of the entire study sample. By repeating this process for both $a:=1$ and $a:=0$, we can then estimate $\psi$ by taking the difference.
Now consider the equivalent set of estimating functions for these discrete steps,
@@
where $\theta=(\beta,\mu^{1},\mu^{0},\psi)$. Here, the first estimating function is for the outcome model restricted to those who completed follow-up, the second and third are for the means of the pseudo-outcomes, and the final estimating function is for the average causal effect. Note that in the first estimating function, $m(\mathbb{X})=E[Y\mid A,X;\beta]$ with the design matrix $\mathbb{X}$ (e.g., $\mathbb{X}=(1,A,X)$ with $\beta=(\beta_{0},\beta_{1},\beta_{2})$). In the case of a continuous outcome, $m$ may be linear regression (i.e., $m(\mathbb{X};\beta)=\mathbb{X}\beta^{T}$). If $Y$ is binary, we might instead use a logistic regression model (i.e., $m(\mathbb{X};\beta)=\text{expit}(\mathbb{X}\beta^{T})$ where expit is the inverse logit function). While these estimating functions may seem to have magically appeared, the first estimating function is simply the score function for the corresponding regression model [15]. As the score functions for many commonly used regression models are well-known results, we simply use those results directly. A similar case holds for the estimating equation for the mean, where $\hat{\mu}^{1}=n^{-1}\sum_{i=1}^{n}\hat{Y}^{1}$ can be rewritten following some algebra as $\sum_{i=1}^{n}\left(\hat{Y}^{1}-\hat{\mu}^{1}\right)=0$. A similar process applies for the third function. The final estimating function does not depend on the data and follows immediately from $\hat{\psi}=\hat{\mu}^{1}-\hat{\mu}^{0}$. For a more detailed step-by-step derivation of these estimating functions for g-computation see the following reference [16].
@@
Having determined the estimating functions for g-computation, we can now show that our estimator is consistent and asymptotically normal. To demonstrate this, M-estimation theory (under some additional regularity conditions) requires us to only show that each of the corresponding estimating equations is unbiased at the true parameter value [15, 14]. So, we need to show $E[g_{1}(Z;\theta)]=0$ holds for the true value of $\theta$. Note that $E\left\{S_{i}\left[Y_{i}-m(\mathbb{X}_{i};\beta)\right]\mathbb{X}_{i}^{T}\right\}=0$ follows directly from standard maximum likelihood theory. For the means, unbiasedness follows from the identification assumptions and correct outcome model specification. Finally, the estimating equation for the average casual effect is unbiased as it is simply a transformation of unbiased parameters. In the Appendix, we provide a more formal proof of these results. Therefore, the g-computation estimator is consistent and asymptotically normal, and we can use the empirical sandwich variance estimator. Note that use of g-computation is still premised on the identification assumptions (e.g., causal consistency, exchangeability, positivity) and that our outcome model is correctly specified.
@@
While the standard g-computation algorithm is valid for some causal structures with selection bias (e.g., traditional differential loss to follow-up where selection, or censoring, depends on the treatment and certain covariates that also impact the outcome), the previous procedure might not work for other structures. Below we consider two case studies where standard g-computation fails to recover the average causal effect but a modified g-computation algorithm does.
@@ 4 Case Study 1: Treatment-induced selection
Consider the example described in Breskin et al. and shown in Figure 1 using a Single World Intervention Graph (SWIG) [13]. Let $A$ denote vaccination status ($1$: vaccine, 00: placebo) and $Y$ denote disease at one year ($1$: yes, 00: no). Here, the vaccine is more likely to result in injection site pain ($X=1$ if pain, $X=0$ otherwise). Additionally, poor overall health ($U$), an unmeasured variable, is related to both the outcome and injection site pain. Finally, $S$ denotes if a person completed follow-up. As such, $Y$ is only observed for those with $S=1$, while $A,X$ are fully observed. The parameter of interest is the average causal effect in the entire trial population, $\psi$. Here, standard g-computation that accounts for $X$ will fail since $X$ is a collider. However, not accounting for $X$ means there is selection bias due to restricting to one level of $S$ which is a descendant of $X$. Therefore, a revised identification strategy is needed. Instead, the interest parameter can be expressed as
@@
for $a\in\{0,1\}$. An identification proof of this result is provided in Appendix 2. This result can how be used to construct a corresponding g-computation estimator.
@@ 4.1 Building an estimator
The result in 2 suggests the following recipe. Starting with the inner expectation, note that $E(Y\mid A,X,S=1)$ can be estimated with a model using only the complete cases. here, $X$ is allowed to be continuous, so we use a parametric model with the parameters $\beta$ (which requires that we assume our model is correctly specified). This model can then be used to generate pseudo-outcomes under $a$ for all observations. The outer expectation indicates that we want to take the mean of the pseudo-outcomes only among those with $A=a$ rather than the entire population. Note that this last step is the key difference from the standard g-computation algorithm given by 1.
This step-by-step process can be translated into the following set of stacked estimating equations,
@@
where $\theta=(\beta,\mu^{1},\mu^{0},\psi)$. This estimating function shares similarities to $g_{1}$. Like $g_{1}$, the outcome model is restricted to those who completed follow-up (i.e., $S=1$). However, the estimating functions for the mean are now conditional on values of $A$. These estimating function can be found by re-expressing the conditional mean as an equation equal to zero. The final estimating function is the average causal effect and is the same as before.
Having determine the estimating equations, we now argue that our proposed estimator is consistent and asymptotically normal. Note that the outcome model is unbiased again following directly from standard maximum likelihood theory. Similarly, the unbiasedness of the conditional means follows from the identification result and correct outcome model specification. Finally, the estimating equation for the average causal effect is unbiased by the definitions of $\psi,\mu^{1},\mu^{0}$. A more formal proof is provided in Appendix 2.
@@ 4.2 Simulation
To illustrate the performance of the proposed estimator, a modified version of the data generating mechanism from Breskin et al. was created (see Appendix 2 for details). In the simulations, we compare taking the difference between the means of the outcomes among complete-cases stratified by treatment arm (naïve analysis), standard g-computation, and our modified g-computation algorithm. For evaluation metrics, we considered bias, empirical standard error (ESE), standard error ratio (SER), and 95% confidence interval (CI) coverage [28]. Bias was defined as the mean of the difference between estimates and the true $\psi$. The ESE was defined as the standard deviation of the point estimates. The SER was defined as the average of the estimated standard errors divided by the ESE. As SER near one indicates the variance is being estimated appropriately, while values below one indicate the variance is being underestimated. CI coverage was defined as the proportion of 95% CIs that covered the true $\psi$. Simulations consisted of $n:=1000$ with 5000 repetitions and were conducted using Python 3.9.4 (Beaveron, OR, USA) using the following packages: NumPy [29], SciPy [30], pandas [31], and delicatessen [23]. We also replicate the results for a single simulated data set in R 4.4.1 (Vienna, Austria).
Results for the simulation study are shown in Table 1. The naïve analysis was biased, as expected. Interestingly, standard g-computation resulted in greater bias relative to the naïve analysis that ignored selection bias for this data generating mechanism. Finally, the modified g-computation estimator had nearly zero bias and the empirical sandwich variance estimator performed well (as indicated by the SER and nominal 95% CI coverage), as was expected.
@@
Bias ESE SER Coverage
Complete-case analysis (naïve) -0.028 0.035 1.00 87%
Standard g-computation -0.144 0.035 1.00 2%
Modified g-computation -0.001 0.036 1.01 95%
@@
ESE: empirical standard error, SER: standard error ratio.

### v2

@@ Abstract
G-computation is a useful estimation method that can be adapted to address various biases in epidemiology. However, these adaptations may not be obvious for some complex causal structures. This challenge is an example of the much wider issue of translating a causal diagram into a novel estimation strategy. To highlight these challenges, we consider two recent cases from the selection bias literature: treatment-induced selection and co-occurrence of biases that lack a joint adjustment set. For each case study, we show how g-computation can be adapted, described how to implement that adaptation, show some general statistical properties, and illustrate the estimator using simulation. To simplify both the theoretical study and practical application of our estimators, we express the proposed g-computation estimators as stacked estimating equations. These examples illustrate how epidemiologists can translate identification results into a g-computation estimator and study the theoretical and finite-sample properties of a novel estimator.
@@ 1 Introduction
A practical challenge to empirical epidemiologic research is translating from a causal diagram identification result to an estimator. While the epidemiologic and statistical literature is filled with novel estimators, these approaches are often tailored to specific scenarios. For example, an estimator may not deal with multiple sources of bias (e.g., confounding, selection bias, missing data), or assume there is an overall sufficient adjustment set for all the sources of systematic errors [1]. The gap between drawing the causal diagram and developing an estimation strategy may dissuade some researchers from considering these novel estimators, even when they are thought to produce more reliable estimates relative to traditional regression methods.
@@
Here, we consider methods to address selection bias as a case study of translating from identification results to estimation. Recent work has focused on clarifying the definition of selection bias [2, 3, 4], and provided graphical rules for addressing selection bias [5, 6]. However, these studies do not immediately provide corresponding estimators to recover treatment effects from selection bias. When faced with selection bias, epidemiologists may default to inverse probability weighting (IPW) estimators, as they allow for one to independently model different sources of bias (e.g., inverse probability of treatment weights for confounding, inverse probability of censoring or sampling weights for different types of selection bias) [7]. While IPW provides a fairly general recipe, there are can be challenges to its application [8]. First, an epidemiologist must correctly specify each IPW model, which can be difficult as it requires sufficient but, oftentimes, distinct background knowledge for each source of bias. Second, the order in which IPW models are constructed can differ depending on the context [9]. Third, IPW estimators are generally less efficient (i.e., wider confidence intervals) relative to competing estimators [10]. These issues can be avoided by adopting outcome-model-based approaches (e.g., g-computation [11]), which only require correct specification of a model for the outcome process (avoiding separate models for each source of bias and concern over the ordering of estimation of different weights) and is generally more efficient [8, 12]. However, g-computation may need to be adapted for certain selection bias mechanisms or under the joint occurrences of biases (e.g., confounding and selection bias) that lack a sufficient adjustment superset [1, 13]. Therefore, adaptations of the standard g-computation algorithm are needed for causal structures under these scenarios.
@@
Using two recent examples in the selection bias literature as case studies, we illustrate how epidemiologists can translate identification results into novel g-computation estimators. For completeness, we rederive the motivating identification results, propose estimators based on those identification results, prove some general properties of our estimators, and illustrate their application using Monte Carlo simulation experiments. For the first case study, we consider treatment-induced selection bias. In the second case study, we study the joint occurrence of confounding and selection bias where there is not an overlapping adjustment set. While estimators are developed for these specific cases, the central ideas of this work can be used to build g-computation estimators in other contexts.
@@ 2 Estimating Equations
M-estimators offer four benefits for us. First, M-estimators provide a general recipe for estimating the variance of parameters via the so-called empirical sandwich variance estimator [15, 19]. By jointly solving a stack of estimating equations, the empirical sandwich variance estimator appropriately incorporates the dependence of parameters on other parameters. Essentially, we can estimate the variance of the parameter of interest, while correctly incorporating the uncertainty in estimation of the outcome model for g-computation. This feature allows us to avoid more computationally intensive methods for variance estimation, such as the bootstrap [20, 21]. However, any variable selection procedure not based on domain knowledge needs to be part of the stacked estimating equations for valid variance esitmation. Note that the empirical sandwich variance estimator can also be used with IPW estimators [22]. Second, M-estimation theory can simplify consistency and asymptotic normality proofs for estimators. Hence, we can more easily prove certain statistical guarantees for our estimator as the sample size approaches infinity. Briefly, if the estimating equations are shown to be unbiased at the true $\theta$, it suffices to show that the corresponding estimator is consistent and asymptotically normal following some additional regularity conditions [15]. This underlying statistical theory is what justifies use of the empirical sandwich variance estimator. As we propose novel estimators, we use this feature to justify our estimators. Third, many of the estimators that epidemiologists routinely use can be readily expressed with estimating equations. For example, both the arithmetic mean and generalized linear model can be expressed as estimating equations [8, 15]. Therefore, the underlying pieces of the estimators proposed in this paper are connected to approaches already used by epidemiologists. Finally, there is free and open-source software for the automated computation of M-estimators [23, 24], simplifying their application.
@@ 3 Standard G-computation
Before the case studies, we review a standard g-computation implementation for the average causal effect with selection bias and how it can be implemented as an M-estimator. Let $Y$ denote the outcome, $A$ denote the treatment, $S$ denote whether a person completed follow-up (i.e., $Y$ is only observed for those with $S=1$), and $X$ be a set of baseline covariates. Following the Single-World Intervention Graph (SWIG) in Figure 1 [13], $X$ is a sufficient adjustment set for both confounding and selection bias due to differential loss to follow-up. The parameter of interest is $\psi=\mu^{1}-\mu^{0}$, where $\mu^{a}=E[Y^{a}]$, $Y^{a}$ is the potential outcome under treatment $a$, and $E[\cdot]$ is the expected value function. $\psi$ can be re-expressed in terms of the observed data as
@@
following conditional exchangeability with positivity for both confounding and selection bias, and causal consistency (see Appendix 1 for details) [25, 26, 27].
Figure 1: Single world intervention graph for confounding and selection bias with a sufficient adjustment superset
Using the expression in 1 and the algorithm described by Snowden et al. adapted for missing outcome [11], we can estimate $\psi$. First, we fit an outcome model for $Y$ given $A,X$ among those who completed follow-up (i.e., $S=1$). Using that model, we then generate predict values of $Y$ under $a$ for all observations in the study sample (both $S=1$ and $S=0$), which we refer to as pseudo-outcomes and denote as $\hat{Y}^{a}$. Then we take the mean of the pseudo-outcomes, $\hat{\mu}^{a}=n^{-1}\sum_{i=1}^{n}\hat{Y}^{a}_{i}$, where $n$ is the size of the entire study sample. By repeating this process for $a:=1$ and $a:=0$, we can then estimate $\psi$ by taking the difference.
Now consider the estimating functions that correspond to these discrete steps,
@@
where $\theta=(\beta,\mu^{1},\mu^{0},\psi)$. Here, the first estimating function is for the outcome model restricted to those who completed follow-up, the second and third are for the means of the pseudo-outcomes, and the final estimating function is for the average causal effect. Note that in the first estimating function, $m(\mathbb{X})=E[Y\mid A,X;\beta]$ with the design matrix $\mathbb{X}$ (e.g., $\mathbb{X}=(1,A,X)$ with $\beta=(\beta_{0},\beta_{1},\beta_{2})$). In the case of a continuous outcome, $m$ may be linear regression (i.e., $m(\mathbb{X};\beta)=\mathbb{X}\beta^{T}$). If $Y$ is binary, we might instead use a logistic regression model (i.e., $m(\mathbb{X};\beta)=\text{expit}(\mathbb{X}\beta^{T})$ where expit is the inverse logit function). While these estimating functions may seem to have magically appeared, the first estimating function is simply the score function for the corresponding regression model [15]. One could derive this estimation equation by taking the derivative of the log-likelihood for the chosen regression model [16]. As the score functions for many commonly used regression models are well-known results, we simply use those results directly. A similar case holds for the estimating equation for the mean, where $\hat{\mu}^{1}=n^{-1}\sum_{i=1}^{n}\hat{Y}^{1}$ can be rewritten following some algebra as $\sum_{i=1}^{n}\left(\hat{Y}^{1}-\hat{\mu}^{1}\right)=0$. A similar process applies for the third function. The final estimating function does not depend on the data and follows immediately from $\hat{\psi}=\hat{\mu}^{1}-\hat{\mu}^{0}$. For a more detailed step-by-step derivation of estimating functions for g-computation see the following reference [16].
@@
Having determined the estimating functions, we can now show that our estimator is consistent and asymptotically normal. To demonstrate this, M-estimation theory (under some additional regularity conditions) requires us to only show that each of the corresponding estimating equations is unbiased at the true parameter value [15, 14]. So, we need to show $E[g_{1}(Z;\theta)]=0$ holds for the true value of $\theta$. Here, we provide a sketch of an argument with a more formal proof in the Appendix. Note that $E\left\{S_{i}\left[Y_{i}-m(\mathbb{X}_{i};\beta)\right]\mathbb{X}_{i}^{T}\right\}=0$ follows directly from standard maximum likelihood theory. For the means, unbiasedness follows from the identification assumptions and correct outcome model specification. Finally, the estimating equation for the average casual effect is unbiased as it is simply a transformation of unbiased parameters. Therefore, the g-computation estimator is consistent and asymptotically normal. Note that use of g-computation is still premised on the identification assumptions (e.g., causal consistency, exchangeability, positivity) and that the outcome model is correctly specified.
@@
While the standard g-computation algorithm is valid for some causal structures with selection bias, the previous procedure might not work for other structures. Below we consider two case studies where standard g-computation fails to recover the average causal effect but a modified g-computation algorithm does.
@@ 4 Case Study 1: Treatment-induced selection
Consider the example described in Breskin et al. and shown in Figure 2 [13]. Let $A$ denote vaccination status (111: vaccine, 0: placebo) and $Y$ denote disease at one year (111: yes, 0: no). Here, the vaccine is more likely to result in injection site pain ($X=1$ if pain, $X=0$ otherwise). Additionally, poor overall health ($U$), an unmeasured variable, is related to both the outcome and injection site pain. Finally, $S$ denotes if a person completed follow-up. As such, $Y$ is only observed for those with $S=1$, while $A,X$ are fully observed. The parameter of interest is the average causal effect in the entire trial population, $\psi$. Here, $U$ only affects selection through $X$ but standard g-computation that incorporates $X$ will be biased since $X$ is a collider. However, not accounting for $X$ means there is selection bias due to an open backdoor path between $S^{a}$ and $Y^{a}$ (type 2 selection bias) [2]. Therefore, a revised identification strategy is needed. Instead, the interest parameter can be expressed as
@@
for $a\in\{0,1\}$ (proof in Appendix 2).This result can how be used to construct a modified g-computation estimator.
@@ 4.1 Building an estimator
The result in 2 suggests the following recipe. Starting with the inner expectation, note that $E(Y\mid A,X,S=1)$ can be estimated with a model using only the complete cases. here, $X$ is allowed to be continuous, so we use a parametric model with the parameters $\beta$ (which requires that we assume our model is correctly specified). This model can then be used to generate pseudo-outcomes under $a$ for all observations. The outer expectation indicates that we want to take the mean of the pseudo-outcomes only among those with $A=a$ rather than the entire population. Note that this last step is the key difference from the standard g-computation algorithm.
This step-by-step process can be translated into the following set of stacked estimating functions,
@@
where $\theta=(\beta,\mu^{1},\mu^{0},\psi)$. This estimating function shares similarities to $g_{1}$. Like $g_{1}$, the outcome model is restricted to those who completed follow-up (i.e., $S=1$). However, the estimating functions for the mean are now conditional on values of $A$. These estimating function can be found by re-expressing the formula for the conditional mean as an equation equal to zero. The final estimating function is the average causal effect.
Having determine the estimating equations, we now argue that our proposed estimator is consistent and asymptotically normal with a more formal proof provided in Appendix 2. Note that the outcome model is unbiased again following directly from standard maximum likelihood theory. Similarly, the unbiasedness of the conditional means follows from the identification result in 2 and correct outcome model specification. Finally, the estimating equation for $\psi$ is unbiased by the definitions of the corresponding parameters.
@@ 4.2 Simulation
To verify our mathematical derivation and illustrate the performance of the proposed estimator, a modified version of the data generating mechanism from Breskin et al. was created (see Appendix 2 for details). In the simulations, we compare taking the difference between the means of the outcomes among complete-cases stratified by treatment arm (naïve analysis), standard g-computation with $X$ (as in 1), and our modified g-computation algorithm. For reference, we also include results for an IPW estimator (Appendix 2). For evaluation metrics, we considered bias, empirical standard error (ESE), root mean squared error (RMSE), standard error ratio (SER), and 95% confidence interval (CI) coverage [28]. Bias was defined as the mean of the difference between estimates and the true $\psi$. ESE was defined as the standard deviation of the point estimates. RMSE was defined as the square root of the mean of the squared difference between estimates and the true $\psi$, where smaller values are better. SER was defined as the average of the estimated standard errors divided by the ESE. As SER near one indicates the variance is being estimated appropriately, while values below one indicate the variance is being underestimated. CI coverage was defined as the proportion of 95% CIs that covered the true $\psi$. Simulations consisted of $n:=1000$ with 5000 repetitions and were conducted using Python 3.9.4 (Beaveron, OR, USA) using the following packages: NumPy [29], SciPy [30], pandas [31], and delicatessen [23]. We also replicate the results for a single simulated data set in R 4.4.1 (Vienna, Austria).
Results for the simulation study are shown in Table 1. The naïve analysis was biased, as expected. For our specific mechanism, standard g-computation resulted in greater bias relative to the naïve analysis. Finally, the modified g-computation and IPW estimators had nearly zero bias and the empirical sandwich variance estimator performed well (as indicated by the SER and nominal 95% CI coverage), as was expected. In terms of the RMSE, the modified g-computation estimator was the best performing estimator.
@@
Bias ESE RMSE SER Coverage
Complete-case analysis (naïve) -0.028 0.035 0.045 1.00 87%
Standard g-computation -0.144 0.035 0.148 1.00 2%
Modified g-computation -0.001 0.036 0.036 1.01 95%
Inverse probability weighting -0.001 0.039 0.039 1.00 95%
@@
ESE: empirical standard error, RMSE: root mean squared error, SER: standard error ratio, CI: confidence interval.